OpenStudy (anonymous):

How do I calculate the molalities of 1.22M sugar (C12H22O11) solution density = 1.12g/mL (1.74m)

OpenStudy (frostbite):

Hey @rgermina5873 and welcome to \(\LARGE \sf \bbox[#40B9E9]{\color{white}{Open}}\bbox[#A8CE91]{\color{white}{Study!}}\) The molality \(b\) is defined as:\[\large b=\frac{ n }{ m }\] We got a concentration and density defined as:\[\large C=\frac{ n }{ V }\] \[\large \rho = \frac{ m }{ V }\] Lets see what happens when we multiply with the concentration with the reciprocal density \(\rho^{-1}\): \[\large C \times \rho^{-1}=\frac{ n }{ V } \times \frac{ V }{ m }=\frac{ n }{ m }=b\] So we see that: \[\large b=\frac{ C }{ \rho }\] Make sure that the units get correct as molality has the unit \([\large \frac{mol}{kg}]\)

OpenStudy (anonymous):

You are good. I could never understand this problem but you just get it. Thanks.

OpenStudy (frostbite):

Thank you, but you understand how we derived our wished equation? :)

OpenStudy (anonymous):

No not really. Please explain

OpenStudy (frostbite):

Okay I try again. molality \(b\) is a measure of how many moles of molecules we got per kilogram. The equation would therefor be: \[\large b=\frac{ n }{ m }\]Where \(n\) is the amount of substance [mol] and m the mass (of the solvent) [kg]. The unit for molality therefor become \(\large \frac{mol}{g}\) Understand so far? :)

OpenStudy (anonymous):

Yes I understand so far.

OpenStudy (frostbite):

Good we now look at the density and concentration: The density \(\rho\) (the sign is called "rho") is defined as a measure how much a substance weights per volume. The equation become: \[\large \rho = \frac{ m }{ V }\] Where \(m\) is the mass [g] and \(V\) the volume [ml]. The concentration \(C\) is a measure of how many mole of molecules we got per volume. The equation become: \[\large C=\frac{ n }{ V }\] The \(n\) and \(V\) you know from before, but what is special this time is the unit: we define as new united called \(M\) and is equal to \(\frac{mol}{l}\)

OpenStudy (frostbite):

So try multiply: The reciprocal densitry \(\rho^{-1}\) and the concentration \(C\) Note: \[\large \rho ^{-1}=\frac{ 1 }{ \frac{ m }{ V } }=\frac{ V }{ m }\]

OpenStudy (anonymous):

I've never seen that formula before.

OpenStudy (anonymous):

Using this formula will help me get to the answer. Thanks once again.

OpenStudy (kayne):

Great work @Frostbite :)